OpenStudy (anonymous):

x + 20 + 10x = 20 + 9x is the answer to this problem 0?

OpenStudy (texaschic101):

x + 20 + 10x = 20 + 9x 11x + 20 = 20 + 9x 11x - 9x = 20 - 20 2x = 0 x = 0 check.. x + 20 + 10x = 20 + 9x 0 + 20 + 10(0) = 20 + 9(0) 20 = 20 (correct) yes...x = 0